OpenStudy (notebook):

Would appreciate if someone checked my optimization answer for: John, the manager of a store, wishes to add a fenced-in rectangular storage yard of 20,000 square feet, using the building as one side of the yard. Find the minimum amount of fencing that must be used to enclose the remaining 3 sides of the yard. My answer: 400 yards

OpenStudy (sshayer):

let x and y be the sides of rectangle. xy=20,000 y=20000/x \[L=2x+y,=2x+\frac{ 20,000 }{ x }\] where L is the length of fence. \[\frac{ dL }{ dx }=2-\frac{ 20000 }{ x^2 }\] \[\frac{ dL }{ dx }=0,gives\] \[2-\frac{ 20000 }{ x^2 }=0,gives \] \[2x^2=20000,x^2=10000,x=100\] \[\frac{ d^2L }{ dx^2 }=-\frac{ -2 \times 20,000 }{ x^3 }=\frac{ 40,000 }{ x^3 }\] at x=100 \[\frac{ d^2L }{ dx^2 }=\frac{ 40,000 }{ 100^3 }>0 \] Hence L is minimum when x=100 \[y=\frac{ 20000 }{ 100 }=200\] Hence minimum length of fence=2x+y=2*100+200=400 yards so you are correct.

OpenStudy (notebook):

Thank you.

OpenStudy (sshayer):

yw